### v2

# Large Sets of Quasi-Complementary Sequences From Polynomials over Finite Fields and

Abstract
Perfect complementary sequence sets (PCSSs) are widely used in multi-carrier code-division multiple-access (MC-CDMA) communication systems. However, the set size of a PCSS is upper bounded by the number of row sequences of each two-dimensional matrix in the PCSS. Then quasi-complementary sequence sets (QCSSs) were proposed to support more users in MC-CDMA communications. For practical applications, it is desirable to construct an $(M,K,N,\vartheta_{\max})$-QCSS with $M$ as large as possible and $\vartheta_{max}$ as small as possible, where $M$ is the number of matrices with $K$ rows and $N$ columns in the set and $\vartheta_{\max}$ denotes its periodic tolerance. There exists a tradeoff among these parameters. Constructing QCSSs achieving or nearly achieving the known correlation lower bound has been an interesting research topic. Up to now, only a few constructions of asymptotically optimal or near-optimal periodic QCSSs have been reported in the literature. In this paper, based on polynomials over finite fields and Gaussian sums, we construct five new families of asymptotically optimal or near-optimal periodic QCSSs with large set sizes and low periodic tolerances. These families of QCSSs have set size $\Theta(K^{2})$ or $\Theta(K^{3})$ and flock size $K$. To the best of our knowledge, only a small amount of known families of periodic QCSSs with set size $\Theta(K^{2})$ have been constructed and most of other known periodic QCSSs have set sizes much smaller than $\Theta(K^{2})$. Our new constructed periodic QCSSs with set size $\Theta(K^{2})$ and flock size $K$ have the best parameters among all known ones. They have larger set sizes or lower periodic tolerances. The periodic QCSSs with set size $\Theta(K^{3})$ and flock size $K$ constructed in this paper have the largest set size among all known families of asymptotically optimal or near-optimal periodic QCSSs.
Index Terms:
Quasi-complementary sequence sets, polynomials over finite fields, Gaussian sums
I Introduction
A perfect complementary sequence set (PCSS), which is also called mutually orthogonal complementary sequence set, consists of a number of two-dimensional matrices. The nontrivial auto-correlation and cross-correlation sums of the row sequences of the matrices are zero for any nonzero time-shift [21]. PCSSs have found many nice applications including peak-to-average power ration reduction [4], inter-symbol interference channel estimation [20], radar waveform design [18] and multi-carrier code-division multiple-access (MC-CDMA) communication systems [2]. The application of PCSSs has been viewed as a very promising technology for the next generation CDMA [3]. In an MC-CDMA communication system, each user is assigned a two-dimensional matrix and all of its row sequences are transmitted over separate sub-carrier channels simultaneously [2]. Then the number of the row sequences of a matrix in the PCSS and that of sub-carrier channels are equal. However, the set size of a PCSS can not be larger than the number of row sequences of each two-dimensional matrix in it by the correlation lower bound. In order to support more users in MC-CDMA communication systems, some researchers proposed two new kinds of complementary sequence sets which are zero-correlation zone complementary sequence sets (ZCZ-CSSs) [5, 8] and low-correlation zone complementary sequence sets (LCZ-CSS) [13]. An LCZ-CSS (resp. ZCZ-CSS) has low (resp. zero) correlation magnitudes within a zone around the zero-shift. In recent years, many researches have been done on constructing LCZ-CSSs and ZCZ-CSSs.
In 2013, Liu et al. proposed another type of complementary sequence sets, i.e. the quasi-complementary sequence sets (QCSSs), which could have larger set sizes than those of PCSSs [14]. In the literature, there are two kinds of QCSSs including periodic QCSSs and aperiodic QCSSs [14, 15]. In this paper, we mainly study periodic QCSSs. A periodic $(M,K,N,\vartheta_{\max})$-QCSS over an alphabet $A$ is a set consisting of $M$ two-dimensional $K\times N$ matrices such that the maximum value $\vartheta_{\max}$ of nontrivial periodic correlation sums of the row sequences is a small positive number. Each row of the two-dimensional matrices is called a constituent sequence of length $N$ and $K$ is called the flock size. The maximum non-trivial periodic correlation sum $\vartheta_{\max}$ measures the maximum possible amount of the multipath interference and the multiuser interference. For the application of a periodic QCSSs in MC-CDMA communications, it is required that its set size $M$ is as large as possible and $\vartheta_{\max}$ is as small as possible. However, there exists a tradeoff among the parameters $M,K,N$ and $\vartheta_{\max}$. In [14], Liu et al. presented a lower bound for the maximum periodic correlation sum of a periodic $(M,K,N,\vartheta_{\max})$-QCSS as follows:
$\vartheta_{\max}\geq\vartheta_{{\text{opt}}}=KN\sqrt{\frac{M/K-1}{MN-1}}.$ (1)
It is easy to verify that $M\leq K$, which is the upper bound of the size of a periodic PCSS, provided that $\vartheta_{\max}=0$. The tightness factor of the above correlation lower bound for a periodic QCSS is defined by $\rho=\vartheta_{\max}/\vartheta_{{\text{opt}}}$. Note that $\rho\geq 1$. A QCSS is said to be optimal if $\rho=1$ and near-optimal if $1<\rho\leq 2$ [14]. Constructing optimal or near-optimal periodic QCSSs has been an interesting research topic in recent years. However, up to now, no infinite family of optimal QCSSs achieving the lower bound in (1) was reported. Only a few families of asymptotically optimal or near-optimal QCSSs have been constructed. In [14], Liu et al. constructed the first family of asymptotically optimal and the first family of asymptotically near-optimal periodic QCSSs from the Single difference sets and the existing optimal quaternary sequence sets. Using a similar idea, Li et al. constructed a family of asymptotically optimal periodic QCSSs from almost difference sets in [12]. In [9], Li et al. generalized the constructions in [14] and derived new asymptotically optimal periodic QCSSs. In [10, 11], Li et al. presented several constructions of asymptotically optimal QCSSs from additive and multiplicative characters over finite fields. Later, Luo et al. gave three new constructions of asymptotically optimal QCSSs with small alphabet sizes by additive characters over finite fields [16]. Recently, Xiao et al. used additive characters of finite fields to construct two more families of asymptotically optimal periodic QCSSs [22]. The parameters of these known families of periodic QCSSs are summarized in Table I. It is shown that only a small amount of known families of periodic QCSSs with set size $\Theta(K^{2})$ and flock size $K$ have been constructed and most of other known periodic QCSSs have set sizes much smaller than $\Theta(K^{2})$. The purpose of this paper is to use new approaches to construct five new families of asymptotically optimal or near optimal QCSSs with better parameters. The main contents of this paper are as follows:
  1. $\bullet$
    Using a family of quadratic polynomials over finite fields, we construct a family of $(q^{2},q,q-1,q)$-QCSSs with alphabet size $p$ and a family of $(2^{2n},2^{n},2^{n}-1,2^{n})$-QCSSs with alphabet size $2$, where $q=p^{n}$ with $p$ an odd prime and $n$ a positive integer. These two families of QCSSs are asymptotically optimal with respect to the correlation lower bound and have larger set sizes than those in [10, 22] by Table I.
  2. $\bullet$
    A family of $(2^{3n},2^{n},2^{n}-1,2^{n+1})$-QCSSs with alphabet size $2$ is constructed from a special quadrinomial over finite fields with characteristic 2. A family of $(p^{3n},p^{n},p^{n}-1,2p^{n})$-QCSSs with alphabet size $p$ is also constructed from a cubic quadrinomial over finite fields with characteristic $p$ for an odd prime $p$. These two families of QCSSs are asymptotically near-optimal with respect to the correlation lower bound. Note that the set size of each of them is the cube of the flock size. In this sense, these two families of QCSSs have the largest set sizes among all known families of asymptotically optimal or near-optimal periodic QCSSs.
  3. $\bullet$
    By Gaussian sums, we construct a family of $(p^{n}(p^{n}-1),p^{n}-1,p^{n}-1,p^{n})$-QCSSs which have lower periodic tolerance than the known $(p^{n}(p^{n}-1),p^{n}-1,p^{n}-1,p^{n}+1)$-QCSSs in [22], though our QCSSs have larger alphabet size.
According to the parameters of all known periodic QCSSs listed in Table I, our QCSSs have new parameters and very large set sizes. Most of them have very small alphabet size. These new QCSSs may be used in MC-CDMA communication systems to support a large number of users.
Table I: The parameters of known periodic QCSSs.
Set size $M$ Flock size $K$ Length $N$ $\vartheta_{\max}$ Alphabet Size Constraint References
$2^{n}$ $12^{n-1}-$ 1 $12^{n}-$ 1 $\frac{2^{n}+2^{\frac{n}{2}}}{2}$ $4(2^{n}-1)$ $n>1$ [14]
$2^{n}$ $12^{n-1}-$ 1 $22^{n+1}-$ 2 $2^{n}+2^{n/2}$ $4(2^{n}-1)$ $n>1$ [14]
$2^{2n}$ $2^{n-1}-2^{n/2}$ $12^{2n}-$ 1 $\leq w$ $4(2^{n}-1)$ $n\geq 3$ [12]
$p$ $\frac{p-1}{2}$ $p$ $\leq\frac{p+\sqrt{p}}{2}$ $p$ $p\equiv 1\pmod{4}$ is a prime [9]
$p^{n}-2$ $\frac{p^{n}-1}{2}$ $p^{n}-1$ $\leq\frac{p^{n}+4\sqrt{p^{n}}+3}{2}$ $p^{n}-1$ $p$ is an odd prime [10]
$p^{2n}-2$ $p^{n}$ $p^{2n}-1$ $p^{n}(p^{2n}+3)$ $p^{2n}-1$ $p$ is a prime [10]
$p^{n}-1$ $\frac{p^{n}-1}{2}$ $p^{n}-1$ $\leq\frac{p^{n}+\sqrt{p^{n}}}{2}$ $p(p^{n}-1)$ $p$ is an odd prime [11]
$p^{n}-1$ $p^{n-1}$ $p^{n}-1$ $\leq p^{n-\frac{1}{2}}$ $p(p^{n}-1)$ $p$ is a prime and $n>1$ [11]
$p^{2n}-1$ $p^{n}$ $p^{2n}-1$ $p^{3n/2}$ $p(p^{2n}-1)$ $p$ is a prime [11]
$12^{n}-$ 1 $12^{n-1}-$ 1 $12^{n}-$ 1 $2^{n-1}$ $2(2^{n}-1)$ $n>1$ [11]
$12^{n}-$ 1 $2^{n-2}$ $12^{n}-$ 1 $\leq 3\cdot 2^{n-2}$ $2(2^{n}-1)$ $n>1$ [11]
$p^{n}$ $\frac{p^{n}-1}{2}$ $p^{n}-1$ $\frac{p^{n}+1}{2}$ $p$ $p$ is an odd prime, $p^{n}>3$, [16]
$p^{n}$ $\frac{p^{n}-p^{n-1}}{2}$ $p^{n}-1$ $\frac{p^{n}+p^{n-1}}{2}$ $p$ $p$ is an odd prime, $p^{n}>3$, [16]
$p^{n}$ $u$ $p^{n}-1$ $v$ $p$ $p$ is an odd prime, $n>1$ is odd, [16]
$p^{2n}-p^{n}$ $p^{n}$ $p^{n}-1$ $p^{n}$ $p$ $p$ is a prime, $n>1$ [22]
$p^{2n}-p^{n}$ $p^{n}-1$ $p^{n}-1$ $p^{n}+1$ $p$ $p$ is a prime, $n>1$ [22]
$p^{2n}$ $p^{n}$ $p^{n}-1$ $p^{n}$ $p$ $p$ is an odd prime, $n\geq 1$ Theorem III.1
$2^{2n}$ $2^{n}$ $12^{n}-$ 1 $2^{n}$ $2$ $n\geq 1$ Theorem III.2
$2^{3n}$ $2^{n}$ $12^{n}-$ 1 $2^{n+1}$ $2$ $n\geq 3$ Theorem III.3
$p^{3n}$ $p^{n}$ $p^{n}-1$ $2p^{n}$ $p$ $p$ is an odd prime, $n\geq 1$, Theorem III.4
$p^{2n}-p^{n}$ $p^{n}-1$ $p^{n}-1$ $p^{n}$ $p(p^{n}-1)$ $p$ is a prime Theorem III.5
  • $u=\frac{p^{n}-p^{n-1}+(-1)^{(p-1)(n+3)/4}(p-1)p^{(n-1)/2}}{2}$, $v=\frac{p^{n}+p^{n-1}-(-1)^{(p-1)(n+3)/4}(p-1)p^{(n-1)/2}}{2}$,
    $w=(1+2^{n})\sqrt{2^{n}+2^{n-2}-2^{n/2}}$
II Preliminaries
In this section, we recall some basic definitions of QCSSs and results on characters over finite fields, Gaussian sums, character sums, and the numbers of zeros of some polynomials over finite fields.
II-A Periodic quasi-complementary sequence set
For two complex-valued sequences $\textbf{a}=(a_{0},a_{1},\cdots,a_{N-1})$ and $\textbf{b}=(b_{0},b_{1},\cdots,b_{N-1})$ of period $N$, the periodic correlation function between them is defined by
$R_{\textbf{a},\textbf{b}}(\tau)=\sum_{t=0}^{N-1}a_{t}\overline{b_{t+\tau}}$
for $10\leq\tau\leq N-$ 1, where $(t+\tau)$ is calculated modulo $N$ and $\overline{b_{t+\tau}}$ is the complex conjugation of $b_{t+\tau}$.
Let =$\mathbb{C}=\left\{\mathbf{C}^{0},\mathbf{C}^{1},\cdots,\mathbf{C}^{M-1}\right\}$ be a set of $M$ complementary sequences such that each $\mathbf{C}^{m}$ is a two-dimensional matrix with size $K\times N$ given by
$\mathbf{C}^{m}=\left[\begin{array}[]{cccc}\mathbf{c}_{0}^{m}\\ \mathbf{c}_{1}^{m}\\ \vdots\\ \mathbf{c}_{K-1}^{m}\\ \end{array}\right]=\left[\begin{array}[]{cccc}c_{0,0}^{m}&c_{0,1}^{m}&\cdots&c_{0,N-1}^{m}\\ c_{1,0}^{m}&c_{1,1}^{m}&\cdots&c_{1,N-1}^{m}\\ \vdots&\vdots&\ddots&\vdots\\ c_{K-1,0}^{m}&c_{K-1,1}^{m}&\cdots&c_{K-1,N-1}^{m}\\ \end{array}\right],$
where $\mathbf{c}_{k}^{m}=(c_{k,0}^{m},c_{k,1}^{m},\cdots,c_{k,N-1}^{m})$ is the $k$-th constituent sequence of length $N$, $10\leq k\leq K-$ 1, $10\leq m\leq M-$ 1. The periodic correlation function between two complementary sequences $\mathbf{C}^{{m}_{1}}$ and $\mathbf{C}^{{m}_{2}}$ is defined by
$R_{{{\textbf{C}}^{{m}_{1}}},{{\textbf{C}}^{{m}_{2}}}}(\tau)=\sum_{k=0}^{K-1}R_{{\mathbf{c}_{k}^{{m}_{1}}},{\mathbf{c}_{k}^{{m}_{2}}}}(\tau),\leavevmode\nobreak\ 0\leq\tau<N,$
where $10\leq m_{1},m_{2}\leq M-$ 1. The maximum periodic auto-correlation magnitude and the maximum periodic cross-correlation magnitude of $\mathbb{C}$ are respectively defined by
$\vartheta_{a}=\max\left\{|R_{{{\textbf{C}}^{{m}}},{{\textbf{C}}^{{m}}}}(\tau)|:0\leq m<M,\ 0<\tau<N\right\},$
$\vartheta_{c}=\max\left\{|R_{{{\textbf{C}}^{{m}_{1}}},{{\textbf{C}}^{{m}_{2}}}}(\tau)|:0\leq m_{1}\neq m_{2}<M,\ 0\leq\tau<N\right\}.$
Define the maximum periodic correlation magnitude (also called periodic tolerance) by
$\vartheta_{\max}=\max\{\vartheta_{c},\vartheta_{a}\}.$
It is an important performance measure of the complementary sequence set $\mathbb{C}$ in practice applications. If $\vartheta_{\max}>0$, then $\mathbb{C}$ is called a periodic $(M,K,N,\vartheta_{\max})$ quasi-complementary sequence set (QCSS). Otherwise, $\mathbb{C}$ is called a periodic perfect complementary sequence set (PCSS). Particularly, a PCSS reduces to a matrix consisting of two row sequences provided that $M=1$ and $K=2$. This matrix is called a Golay complementary pair.
II-B Additive and multiplicative characters of finite fields
For a prime $p$ and a positive integer $n$, let $q=p^{n}$ and $\zeta_{p}$ be the primitive $p$-th root of complex unity. Let ${\mathbb{F}}_{q}$ denote the finite field with $q$ elements. Define the trace function from ${\mathbb{F}}_{q}$ onto ${\mathbb{F}}_{p}$ by
${\mathrm{Tr}}_{q/p}(x)=x+x^{p}+\cdots+x^{p^{n-1}},\ x\in{\mathbb{F}}_{q}.$
An additive character of ${\mathbb{F}}_{q}$ is defined as a homomorphism $\chi$ from ${\mathbb{F}}_{q}$ to $\mathbb{C}^{*}$ such that $\chi(x+y)=\chi(x)\chi(y)$ for any $x,y\in{\mathbb{F}}_{q}$. Then
$\chi_{a}(x)=\zeta_{p}^{{\mathrm{Tr}}_{q/p}(ax)},\ x\in{\mathbb{F}}_{q},$
defines an additive character for each $a\in{\mathbb{F}}_{q}$. By definition, $\chi_{a}(x)=\chi_{1}(ax)$. $\chi_{0}$ and $\chi_{1}$ are called the trivial and canonical additive characters, respectively. The complex conjugate $\overline{\chi_{a}}$ of $\chi_{a}$ is defined by $\overline{\chi_{a}}(x)=\overline{\chi_{a}(x)}=\chi_{a}(-x)$ for $x\in{\mathbb{F}}_{q}$. From [17], the orthogonal relation of additive characters is given by
$\sum_{x\in{\mathbb{F}}_{q}}\chi_{1}(ax)=\left\{\begin{array}[]{ll}q&\mbox{if $a=0$},\\ 0&\mbox{if $a\in{\mathbb{F}}_{q}^{*}$}.\end{array}\right.$
Let $\alpha$ be a primitive element of ${\mathbb{F}}_{q}$ and $\zeta_{q-1}$ be the primitive $(q-1)$-th root of complex unity. A multiplicative character of ${\mathbb{F}}_{q}$ is defined as a homomorphism $\varphi$ from ${\mathbb{F}}_{q}^{*}$ to $\mathbb{C}^{*}$ such that $\varphi(xy)=\varphi(x)\varphi(y)$ for any $x,y\in{\mathbb{F}}_{q}^{*}$. Then
$\varphi_{j}(\alpha^{k})=\zeta_{q-1}^{jk},\ k=0,1,\cdots,q-2,$
for each $j=0,1,\cdots,q-2$, defines a multiplicative character. In particular, $\varphi_{0}$ is called the trivial multiplicative character. For odd $q$, if $j=\frac{q-1}{2}$, then $\eta:=\varphi_{\frac{q-1}{2}}$ is called the quadratic multiplicative character of ${\mathbb{F}}_{q}$. The complex conjugate $\overline{\varphi}$ of a multiplicative character $\varphi$ is defined by $\overline{\varphi}(x)=\overline{\varphi(x)}=\varphi(x^{-1})$ for $x\in{\mathbb{F}}_{q}^{*}$. From [17], the orthogonal relation of multiplicative characters is given by
$\sum_{x\in{\mathbb{F}}_{q}^{*}}\varphi_{j}(x)=\left\{\begin{array}[]{ll}q-1&\mbox{if $j=0$},\\ 0&\mbox{if $1\leq j\leq q-2$}.\end{array}\right.$
II-C Gaussian sums and character sums over finite fields
The Gaussian sum $G(\varphi,\chi)$ for a multiplicative character $\varphi$ and an additive character $\chi$ of ${\mathbb{F}}_{q}$ is defined by
$G(\varphi,\chi)=\sum_{x\in{\mathbb{F}}_{q}^{*}}\varphi(x)\chi(x).$
The values of Gaussian sums are different to determine. In some special cases, the explicit values of Gaussian sums are known as follows.
Lemma II.1.
[17] Let $q=p^{n}$ with odd prime $p$. Then
$G(\eta,\chi_{1})$ $=$ $(-1)^{n-1}(\sqrt{-1})^{(\frac{p-1}{2})^{2}n}\sqrt{q}$
$=$ $\left\{\begin{array}[]{lll}(-1)^{n-1}\sqrt{q}&\mbox{for }p\equiv 1\pmod{4},\\ (-1)^{n-1}(\sqrt{-1})^{n}\sqrt{q}&\mbox{for }p\equiv 3\pmod{4}.\end{array}\right.$
Lemma II.2.
[17] Let $\varphi$ be a multiplicative character of ${\mathbb{F}}_{q}$ and $\chi$ be an additive character of ${\mathbb{F}}_{q}$. Then
$G(\varphi,\chi)$ $=$ $\left\{\begin{array}[]{lll}q-1&\text{if}\ \varphi=\varphi_{0},\chi=\chi_{0},\\ -1&\text{if}\ \varphi=\varphi_{0},\chi\neq\chi_{0},\\ 0&\text{if}\ \varphi\neq\varphi_{0},\chi=\chi_{0}.\end{array}\right.$
If $\varphi\neq\varphi_{0}$ and $\chi\neq\chi_{0}$, then $|G(\varphi,\chi_{1})|=\sqrt{q}$.
The following character sums will be used in this paper.
Lemma II.3.
[17] Let $\chi$ be a nontrivial additive character of ${\mathbb{F}}_{q}$ and $q$ be a power of an odd prime. Let $f(x)=a_{2}x^{2}+a_{1}x+a_{0}\in{\mathbb{F}}_{q}[x]$ with $a_{2}\neq 0$. Then
$\sum_{x\in{\mathbb{F}}_{q}}\chi(f(x))=\chi(a_{0}-a_{1}^{2}(4a_{2})^{-1})\eta(a_{2})G(\eta,\chi).$
II-D The number of zeros of some polynomials over finite fields
In the following, we present some known results on the numbers of zeros of some polynomials over finite fields.
Lemma II.4.
[19] Let ${\mathbb{F}}_{q}$ be a finite field of characteristic $2$ and let $f(x)=ax^{2}+bx+c\in{\mathbb{F}}_{q}[x]$ be a polynomial of degree $2$. Then
  1. 1.
    $f$ has exactly one zero in ${\mathbb{F}}_{q}$ if and only if $b=0$;
  2. 2.
    $f$ has exactly two zeros in ${\mathbb{F}}_{q}$ if and only if $b\neq 0$ and ${\mathrm{Tr}}_{q/2}(\frac{ac}{b^{2}})=0$;
  3. 3.
    $f$ has no zero in ${\mathbb{F}}_{q}$ if and only if $b\neq 0$ and ${\mathrm{Tr}}_{q/2}(\frac{ac}{b^{2}})=1$.
Now we consider the polynomial
$f(x)=x^{p^{k}+1}+ax^{p^{k}}+bx+c\in\mathbb{F}_{q}[x].$
For $a\neq 0$, we substitute $x$ by $x-a$ in $f(x)$. Let $\alpha_{1}:=b-a^{p^{k}}$ and $\beta_{1}:=c-ab$. Then $f(x)$ can be reduced to $g(x)=x^{p^{k}+1}+\alpha_{1}x+\beta_{1}$. If we substitute $x$ by $ux$ in $g(x)$ for $u^{p^{k}}=\alpha$ and $\alpha=b-a^{p^{k}}\neq 0$, then $g(x)$ can be transformed into the form $P_{\gamma}(x)=x^{p^{k}+1}+x+\gamma$, where $\gamma=\frac{\beta}{u^{p^{k}+1}}$. By the results in [1], we can easily derive the number of zeros of $f(x)$ in ${\mathbb{F}}_{q}$.
Lemma II.5.
[23] Let $k$ and $n$ be positive integers with $\gcd(k,n)=h$. Let $f(x)=x^{p^{k}+1}+ax^{p^{k}}+bx+c$, where $a,b,c\in{\mathbb{F}}_{q}$. The number of zeros of $f(x)$ in ${\mathbb{F}}_{q}$ is denoted by $N_{f}$.
  1. 1.
    If $a=0,b=0$ or $a\neq 0,b=a^{p^{k}}$, then
    $N_{f}=\left\{\begin{array}[]{ll}1&\text{if}\leavevmode\nobreak\ \frac{n}{h}\leavevmode\nobreak\ \text{is}\leavevmode\nobreak\ \text{odd}\leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ p=2,\\ 0,1\leavevmode\nobreak\ \text{or}\leavevmode\nobreak\ 2&\text{if}\leavevmode\nobreak\ \frac{n}{h}\leavevmode\nobreak\ \text{is}\leavevmode\nobreak\ \text{odd}\leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ p\leavevmode\nobreak\ \text{is}\leavevmode\nobreak\ \text{odd},\\ 0,1\leavevmode\nobreak\ \text{or}\leavevmode\nobreak\ p^{h}+1&\text{if}\leavevmode\nobreak\ \frac{n}{h}\leavevmode\nobreak\ \text{is}\leavevmode\nobreak\ \text{even}.\end{array}\right.$
  2. 2.
    If $a=0,b\neq 0$ or $a\neq 0,b\neq a^{p^{k}}$, then $N_{f}$ takes either $0,1,2$ or $p^{h}+1$.
Corollary II.6.
Let $h=1$, $n\geq 3$ and $p=2$ in Lemma II.5. Then $N_{f}\in\{0,1,2,3\}$ and there exists $(a,b,c)\in{\mathbb{F}}_{q}^{3}$ such that $N_{f}=3$.
Proof.
By Lemma II.5, we have $N_{f}\in\{0,1,2,3\}$. According to [7, Theorems 8-10] and the relationship between $f(x)$ and $P_{\gamma}(x)$, it is easy to deduce that there exists $(a,b,c)\in{\mathbb{F}}_{q}^{3}$ such that $N_{f}=3$. ∎
Lemma II.7.
[23] Let $q=p^{n}$ and $k$ be a positive integer with $\gcd(k,n)=h$. Then the trinomial $x^{p^{k}}-ax-b$ has either none, one, or $p^{h}$ zeros in ${\mathbb{F}}_{q}$, where $a,b\in{\mathbb{F}}_{q}$.
III New constructions of periodic QCSSs
In this section, we will present several new constructions of asymptotically optimal or near-optimal periodic QCSSs with large set sizes from special polynomials over finite fields or Gaussian sums.
III-A The first construction of periodic QCSSs from quadratic polynomials
Let $q=p^{n}$ for a prime $p$ and a positive integer $n$. Let $\alpha$ be a primitive element of ${\mathbb{F}}_{q}$. Assume that $d_{0},d_{1},\cdots,d_{q-1}$ are all the elements of ${\mathbb{F}}_{q}$, i.e. ${\mathbb{F}}_{q}=\{d_{0},d_{1},\cdots,d_{q-1}\}$. Let $f(x)=x^{2}+ax+b\in{\mathbb{F}}_{q}[x]$. Let $\chi_{1}$ be the canonical additive character of ${\mathbb{F}}_{q}$. We define a constituent sequence $\mathbf{c}_{l}^{a,b}$ of period $q-1$ as
$\mathbf{c}_{l}^{a,b}=\left(\mathbf{c}_{l}^{a,b}(t)\right)_{t=0}^{q-2},\mbox{ where}\ \mathbf{c}_{l}^{a,b}(t)=\chi_{1}(\alpha^{t}f(d_{l}))\mbox{for}\leavevmode\nobreak\ 0\leq l\leq q-1.$
Then we obtain a two-dimensional $q\times(q-1)$ matrix
$\mathbf{C}^{a,b}=\left[\begin{array}[]{cccc}\mathbf{c}_{0}^{a,b}\\ \mathbf{c}_{1}^{a,b}\\ \vdots,\\ \mathbf{c}_{q-1}^{a,b}\end{array}\right].$
Such matrixes yield a complementary sequence set given by
$\mathcal{C}=\left\{\mathbf{C}^{a,b}:a\in{\mathbb{F}}_{q},b\in{\mathbb{F}}_{q}\right\}.$ (9)
Firstly, we study the case that $p$ is an odd prime.
Theorem III.1.
Let $q=p^{n}$ for odd prime $p$ and positive integer $n$. Let $\mathcal{C}$ be the complementary sequence set defined in (9). Then $\mathcal{C}$ is a periodic $(q^{2},q,q-1,q)$-QCSS with alphabet size $p$ which is asymptotically optimal with respect to the lower bound in (1).
Proof.
For any two complementary sequences $\mathbf{C}^{a_{1},b_{1}}$ and $\mathbf{C}^{a_{2},b_{2}}$ of $\mathcal{C}$ and $\tau\in[0,q-2]$, we have
$R_{\mathbf{C}^{a_{1},b_{1}},\mathbf{C}^{a_{2},b_{2}}}(\tau)$
$=$ $\sum_{l=0}^{q-1}R_{\mathbf{c}_{l}^{a_{1},b_{1}},\mathbf{c}_{l}^{a_{2},b_{2}}}(\tau)$
$=$ $\sum_{l=0}^{q-1}\sum_{t=0}^{q-2}\chi_{1}(\alpha^{t}(d_{l}^{2}+a_{1}d_{l}+b_{1}))\overline{\chi_{1}}(\alpha^{t+\tau}(d_{l}^{2}+a_{2}d_{l}+b_{2}))$
$=$ $\sum_{l=0}^{q-1}\sum_{t=0}^{q-2}\chi_{1}((1-\alpha^{\tau})\alpha^{t}d_{l}^{2}+(a_{1}-a_{2}\alpha^{\tau})\alpha^{t}d_{l}+(b_{1}-b_{2}\alpha^{\tau})\alpha^{t}),$
where $a_{1},b_{1},a_{2},b_{2}\in{\mathbb{F}}_{q}$. We now consider the following cases to determine the value distribution of $R_{\mathbf{C}^{a_{1},b_{1}},\mathbf{C}^{a_{2},b_{2}}}(\tau)$.
Case 1: If $\tau=0$, $a_{1}=a_{2}$ and $b_{1}\neq b_{2}$, then
$R_{\mathbf{C}^{a_{1},b_{1}},\mathbf{C}^{a_{2},b_{2}}}(\tau)$
$=$ $\sum_{l=0}^{q-1}\sum_{t=0}^{q-2}\chi_{1}((b_{1}-b_{2})\alpha^{t}).$
By the orthogonal relation of the additive characters, we have
$R_{\mathbf{C}^{a_{1},b_{1}},\mathbf{C}^{a_{2},b_{2}}}(\tau)=-q.$
Case 2: If $\tau=0$, $a_{1}\neq a_{2}$ and $b_{1}=b_{2}$, then
$R_{\mathbf{C}^{a_{1},b_{1}},\mathbf{C}^{a_{2},b_{2}}}(\tau)$
$=$ $\sum_{l=0}^{q-1}\sum_{t=0}^{q-2}\chi_{1}((a_{1}-a_{2})\alpha^{t}d_{l})$
$=$ $\sum_{t=0}^{q-2}\sum_{x\in{\mathbb{F}}_{q}}\chi_{1}((a_{1}-a_{2})\alpha^{t}x).$
By the orthogonal relation of the additive characters, we have
$R_{\mathbf{C}^{a_{1},b_{1}},\mathbf{C}^{a_{2},b_{2}}}(\tau)=0.$
Case 3: If $\tau=0$, $a_{1}\neq a_{2}$ and $b_{1}\neq b_{2}$, then
$R_{\mathbf{C}^{a_{1},b_{1}},\mathbf{C}^{a_{2},b_{2}}}(\tau)$
$=$ $\sum_{l=0}^{q-1}\sum_{t=0}^{q-2}\chi_{1}((a_{1}-a_{2})\alpha^{t}d_{l})\chi_{1}((b_{1}-b_{2})\alpha^{t})$
$=$ $\sum_{t=0}^{q-2}\chi_{1}((b_{1}-b_{2})\alpha^{t})\sum_{x\in{\mathbb{F}}_{q}}\chi_{1}((a_{1}-a_{2})\alpha^{t}x).$
By the orthogonal relation of the additive characters, we have
$R_{\mathbf{C}^{a_{1},b_{1}},\mathbf{C}^{a_{2},b_{2}}}(\tau)=0.$
Case 4: If $\tau\neq 0$, then $1-\alpha^{\tau}\neq 0$. According to Lemma II.3, we have
$R_{\mathbf{C}^{a_{1},b_{1}},\mathbf{C}^{a_{2},b_{2}}}(\tau)$
$=$ $\sum_{t=0}^{q-2}\sum_{x\in{\mathbb{F}}_{q}}\chi_{1}((1-\alpha^{\tau})\alpha^{t}x^{2}+(a_{1}-a_{2}\alpha^{\tau})\alpha^{t}x+(b_{1}-b_{2}\alpha^{\tau})\alpha^{t})$
$=$ $\sum_{t=0}^{q-2}\chi_{1}\left((b_{1}-b_{2}\alpha^{\tau})\alpha^{t}-\frac{((a_{1}-a_{2}\alpha^{\tau})\alpha^{t})^{2}}{4(1-\alpha^{\tau})\alpha^{t}}\right)\eta((1-\alpha^{\tau})\alpha^{t})G(\eta,\chi_{1})$
$=$ $G(\eta,\chi_{1})\sum_{x\in{\mathbb{F}}_{q}^{*}}\chi_{1}\left(\left(b_{1}-b_{2}\alpha^{\tau}-\frac{(a_{1}-a_{2}\alpha^{\tau})^{2}}{4(1-\alpha^{\tau})}\right)x\right)\eta((1-\alpha^{\tau})x).$
If $b_{1}-b_{2}\alpha^{\tau}=\frac{(a_{1}-a_{2}\alpha^{\tau})^{2}}{4(1-\alpha^{\tau})}$, by the orthogonal relation of the multiplicative characters, we have
$R_{\mathbf{C}^{a_{1},b_{1}},\mathbf{C}^{a_{2},b_{2}}}(\tau)=0.$
If $b_{1}-b_{2}\alpha^{\tau}\neq\frac{(a_{1}-a_{2}\alpha^{\tau})^{2}}{4(1-\alpha^{\tau})}$, then
$R_{\mathbf{C}^{a_{1},b_{1}},\mathbf{C}^{a_{2},b_{2}}}(\tau)$
$=$ $G(\eta,\chi_{1})\eta(1-\alpha^{\tau})\eta\left(b_{1}-b_{2}\alpha^{\tau}-\frac{(a_{1}-a_{2}\alpha^{\tau})^{2}}{4(1-\alpha^{\tau})}\right)$
$\cdot\sum_{x\in{\mathbb{F}}_{q}^{*}}\chi_{1}\left(\left(b_{1}-b_{2}\alpha^{\tau}-\frac{(a_{1}-a_{2}\alpha^{\tau})^{2}}{4(1-\alpha^{\tau})}\right)x\right)\eta\left(\left(b_{1}-b_{2}\alpha^{\tau}-\frac{(a_{1}-a_{2}\alpha^{\tau})^{2}}{4(1-\alpha^{\tau})}\right)x\right)$
$=$ $G(\eta,\chi_{1})^{2}\eta(1-\alpha^{\tau})\eta\left(b_{1}-b_{2}\alpha^{\tau}-\frac{(a_{1}-a_{2}\alpha^{\tau})^{2}}{4(1-\alpha^{\tau})}\right),$
where $G(\eta,\chi_{1})=(-1)^{n-1}(\sqrt{-1})^{(\frac{p-1}{2})^{2}n}\sqrt{q}$ by Lemma II.1. Thus, $|R_{\mathbf{C}^{a_{1},b_{1}},\mathbf{C}^{a_{2},b_{2}}}(\tau)|\in\{0,q\}$.
Summarizing the above four cases, we derive that the maximum periodic correlation magnitude of $\mathcal{C}$ is $q$. Next, we will show that the parameters of the obtained periodic QCSS asymptotically achieve the correlation lower bound in (1). Since $\mathcal{C}$ is a periodic $(q^{2},q,q-1,q)$-QCSS, according to the bound in (1), we have
$\vartheta_{{\text{opt}}}=q(q-1)\sqrt{\frac{q-1}{q^{2}(q-1)-1)}}=q\sqrt{\frac{(q-1)^{3}}{q^{2}(q-1)-1)}}.$
It is easy to see that
$\lim_{q\rightarrow+\infty}\frac{\vartheta_{\max}}{\vartheta_{{\text{opt}}}}=\lim_{q\rightarrow+\infty}\frac{q}{q\sqrt{\frac{(q-1)^{3}}{q^{2}(q-1)-1)}}}=1.$
This completes the proof of this theorem.
Example 1.
Let $p=5$ and $n=2$. Then the parameters of the QCSS $\mathcal{C}$ constructed in Theorem III.1 are $(625,25,24,25)$ and its alphabet is given hy $\{e^{2\pi\sqrt{-1}i/5}:i\in[0,4]\}$. By magma program, we obtain that the periodic correlation function $R_{\mathbf{C}^{a_{1},b_{1}},\mathbf{C}^{a_{2},b_{2}}}(\tau)$ is equal to 00 or $25$ for any $a_{1},b_{1},a_{2},b_{2}\in{\mathbb{F}}_{q}$ and $0\leq\tau\leq 23$ except the trivial case that $a_{1}=a_{2},b_{1}=b_{2}$ and $\tau=0$. For instance, the matrices $\mathbf{C}^{1,\alpha^{0}}$ and $\mathbf{C}^{1,\alpha^{6}}$ in $\mathcal{C}$ are presented as follows, where each entry stands for a power of $\zeta_{5}=e^{2\pi\sqrt{-1}/5}$. By Python program, we show the autocorrelation magnitude distribution of $\mathbf{C}^{1,\alpha^{0}}$ in Fig. 1, the autocorrelation magnitude distribution of $\mathbf{C}^{1,\alpha^{6}}$ in Fig. 2, and the correlation magnitude distribution of $\mathbf{C}^{1,\alpha^{0}}$ and $\mathbf{C}^{1,\alpha^{6}}$ in Fig. 3, respectively.
$\mathbf{C}^{1,\alpha^{0}}=\left[\begin{array}[]{cccc}310332120114240223430441\\ 332120114240223430441310\\ 223430441310332120114240\\ 424022343044131033212011\\ 234304413103321201142402\\ 223430441310332120114240\\ 240223430441310332120114\\ 011424022343044131033212\\ 000000000000000000000000\\ 103321201142402234304413\\ 332120114240223430441310\\ 234304413103321201142402\\ 120114240223430441310332\\ 430441310332120114240223\\ 201142402234304413103321\\ 201142402234304413103321\\ 000000000000000000000000\\ 430441310332120114240223\\ 310332120114240223430441\\ 212011424022343044131033\\ 011424022343044131033212\\ 212011424022343044131033\\ 424022343044131033212011\\ 103321201142402234304413\\ 120114240223430441310332\\ \end{array}\right],\ \mathbf{C}^{1,\alpha^{6}}=\left[\begin{array}[]{cccc}430441310332120114240223\\ 402234304413103321201142\\ 343044131033212011424022\\ 044131033212011424022343\\ 304413103321201142402234\\ 343044131033212011424022\\ 310332120114240223430441\\ 131033212011424022343044\\ 120114240223430441310332\\ 223430441310332120114240\\ 402234304413103321201142\\ 304413103321201142402234\\ 240223430441310332120114\\ 000000000000000000000000\\ 321201142402234304413103\\ 321201142402234304413103\\ 120114240223430441310332\\ 000000000000000000000000\\ 430441310332120114240223\\ 332120114240223430441310\\ 131033212011424022343044\\ 332120114240223430441310\\ 044131033212011424022343\\ 223430441310332120114240\\ 240223430441310332120114\\ \end{array}\right].$
Figure 1: The autocorrelation magnitude distribution of $\mathbf{C}^{1,\alpha^{0}}$ in Example 1
Figure 2: The autocorrelation magnitude distribution of $\mathbf{C}^{1,\alpha^{6}}$ in Example 1
Figure 3: The correlation magnitude distribution of $\mathbf{C}^{1,\alpha^{0}}$ and $\mathbf{C}^{1,\alpha^{6}}$ in Example 1
Secondly, we consider the case that $p=2$ in the following.
Theorem III.2.
Let $q=2^{n}$ for a positive integer $n$. Let $\mathcal{C}$ be the complementary sequence set defined in (9). Then $\mathcal{C}$ is a periodic $(q^{2},q,q-1,q)$-QCSS with alphabet size $2$ which is asymptotically optimal with respect to the correlation lower bound in (1).
Proof.
For any two complementary sequences $\mathbf{C}^{a_{1},b_{1}}$, $\mathbf{C}^{a_{2},b_{2}}$ in $\mathcal{C}$ and $\tau\in[0,q-2]$, we have
$R_{\mathbf{C}^{a_{1},b_{1}},\mathbf{C}^{a_{2},b_{2}}}(\tau)$
$=$ $\sum_{l=0}^{q-1}R_{\mathbf{c}_{l}^{a_{1},b_{1}},\mathbf{c}_{l}^{a_{2},b_{2}}}(\tau)$
$=$ $\sum_{l=0}^{q-1}\sum_{t=0}^{q-2}\chi_{1}\left(\alpha^{t}(d_{l}^{2}+a_{1}d_{l}+b_{1})\right)\overline{\chi_{1}}\left(\alpha^{t+\tau}(d_{l}^{2}+a_{2}d_{l}+b_{2})\right)$
$=$ $\sum_{l=0}^{q-1}\sum_{t=0}^{q-2}\chi_{1}\left((1-\alpha^{\tau})\alpha^{t}d_{l}^{2}+(a_{1}-a_{2}\alpha^{\tau})\alpha^{t}d_{l}+(b_{1}-b_{2}\alpha^{\tau})\alpha^{t}\right),$
where $a_{1},b_{1},a_{2},b_{2}\in{\mathbb{F}}_{q}$. Then we consider the following cases to determine the value distribution of $R_{\mathbf{C}^{a_{1},b_{1}},\mathbf{C}^{a_{2},b_{2}}}(\tau)$.
Case 1: If $\tau=0$, similarly to the proof of Theorem III.1, we have
$R_{\mathbf{C}^{a_{1},b_{1}},\mathbf{C}^{a_{2},b_{2}}}(\tau)=\left\{\begin{array}[]{ll}-q&\text{if}\leavevmode\nobreak\ a_{1}=a_{2},b_{1}\neq b_{2},\\ 0&\text{if}\leavevmode\nobreak\ a_{1}\neq a_{2},b_{1}=b_{2},\\ 0&\text{if}\leavevmode\nobreak\ a_{1}\neq a_{2},b_{1}\neq b_{2}.\end{array}\right.$
Case 2: If $\tau\neq 0$ and $a_{1}=a_{2}\alpha^{\tau}$, then
$R_{\mathbf{C}^{a_{1},b_{1}},\mathbf{C}^{a_{2},b_{2}}}(\tau)$
$=$ $\sum_{l=0}^{q-1}\sum_{t=0}^{q-2}\chi_{1}\left((1-\alpha^{\tau})\alpha^{t}d_{l}^{2}+(b_{1}-b_{2}\alpha^{\tau})\alpha^{t}\right)$
$=$ $\sum_{t=0}^{q-2}\sum_{x\in{\mathbb{F}}_{q}}\chi_{1}\left(\left((1-\alpha^{\tau})x^{2}+(b_{1}-b_{2}\alpha^{\tau})\right)\alpha^{t}\right)$
$=$ $\sum_{x\in{\mathbb{F}}_{q}}\sum_{y\in{\mathbb{F}}_{q}^{*}}\chi_{1}\left(\left((1-\alpha^{\tau})x^{2}+(b_{1}-b_{2}\alpha^{\tau})\right)y\right).$
By Lemma II.4, we know that $f(x)=(1-\alpha^{\tau})x^{2}+b_{1}-b_{2}\alpha^{\tau}$ has exactly one zero in ${\mathbb{F}}_{q}$. By the orthogonal relation of the additive characters, we have
$R_{\mathbf{C}^{a_{1},b_{1}},\mathbf{C}^{a_{2},b_{2}}}(\tau)$
$=$ $(q-1)+(q-1)(-1)$
$=$ $0$
Case 3: If $\tau\neq 0$ and $a_{1}\neq a_{2}\alpha^{\tau}$, then
$R_{\mathbf{C}^{a_{1},b_{1}},\mathbf{C}^{a_{2},b_{2}}}(\tau)$
$=$ $\sum_{x\in{\mathbb{F}}_{q}}\sum_{y\in{\mathbb{F}}_{q}^{*}}\chi_{1}\left((1-\alpha^{\tau})yx^{2}+(a_{1}-a_{2}\alpha^{\tau})yx+(b_{1}-b_{2}\alpha^{\tau})y\right)$
$=$ $\sum_{x\in{\mathbb{F}}_{q}}\sum_{y\in{\mathbb{F}}_{q}^{*}}\chi_{1}\left(\left((1-\alpha^{\tau})x^{2}+(a_{1}-a_{2}\alpha^{\tau})x+(b_{1}-b_{2}\alpha^{\tau})\right)y\right).$
By Lemma II.4, we know that $f(x)=(1-\alpha^{\tau})x^{2}+(a_{1}-a_{2}\alpha^{\tau})x+(b_{1}-b_{2}\alpha^{\tau})$ has exactly two zeros if ${\mathrm{Tr}}_{q/2}\left(\frac{(1-\alpha^{\tau})(b_{1}-b_{2}\alpha^{\tau})}{(a_{1}-a_{2}\alpha^{\tau})^{2}}\right)=0$, and no zero if ${\mathrm{Tr}}_{q/2}\left(\frac{(1-\alpha^{\tau})(b_{1}-b_{2}\alpha^{\tau})}{(a_{1}-a_{2}\alpha^{\tau})^{2}}\right)=1$. By the orthogonal relation of the additive characters, we have
$R_{\mathbf{C}^{a_{1},b_{1}},\mathbf{C}^{a_{2},b_{2}}}(\tau)=\left\{\begin{array}[]{ll}-q&\text{if}\leavevmode\nobreak\ {\mathrm{Tr}}_{q/2}\left(\frac{(1-\alpha^{\tau})(b_{1}-b_{2}\alpha^{\tau})}{(a_{1}-a_{2}\alpha^{\tau})^{2}}\right)=1,\\ q&\text{if}\leavevmode\nobreak\ {\mathrm{Tr}}_{q/2}\left(\frac{(1-\alpha^{\tau})(b_{1}-b_{2}\alpha^{\tau})}{(a_{1}-a_{2}\alpha^{\tau})^{2}}\right)=0.\end{array}\right.$
Based on all the cases discussed above, we deduce that the maximum periodic correlation magnitude of $\mathcal{C}$ is $q$. Then $\mathcal{C}$ is a periodic $(q^{2},q,q-1,q)$-QCSS which is asymptotically optimal with respect to the correlation bound in (1) by a similar proof as that of Theorem III.1. This completes the proof of this theorem. ∎
Example 2.
Let $p=2$ and $n=4$. Then the parameters of the QCSS $\mathcal{C}$ constructed in Theorem III.2 are $(256,16,15,16)$ and its alphabet is given by $\{(-1)^{i}:i\in[0,1]\}$. By Magma program, we verify that the periodic correlation function $R_{\mathbf{C}^{a_{1},b_{1}},\mathbf{C}^{a_{2},b_{2}}}(\tau)$ is equal to 00 or $16$ for any $a_{1},b_{1},a_{2},b_{2}\in{\mathbb{F}}_{q}$ and $0\leq\tau\leq 14$ except the trivial case that $a_{1}=a_{2},b_{1}=b_{2}$ and $\tau=0$. For instance, the matrices $\mathbf{C}^{1,\alpha}$ and $\mathbf{C}^{1,\alpha^{2}}$ in $\mathcal{C}$ are presented as follows, where each entry stands for a power of $\zeta_{2}=-1$. By Python program, we show the autocorrelation magnitude distribution of $\mathbf{C}^{1,\alpha}$ in Fig. 4, the autocorrelation magnitude distribution of $\mathbf{C}^{1,\alpha^{2}}$ in Fig. 5, and the correlation magnitude distribution of $\mathbf{C}^{1,\alpha}$ and $\mathbf{C}^{1,\alpha^{2}}$ in Fig. 6, respectively.
$\mathbf{C}^{1,\alpha}=\left[\begin{array}[]{cccc}010011010111100\\ 100110101111000\\ 101111000100110\\ 110101111000100\\ 100110101111000\\ 011010111100010\\ 001001101011110\\ 000000000000000\\ 101111000100110\\ 000000000000000\\ 011010111100010\\ 111100010011010\\ 111100010011010\\ 001001101011110\\ 110101111000100\\ 010011010111100\\ \end{array}\right],\ \mathbf{C}^{1,\alpha^{2}}=\left[\begin{array}[]{cccc}100110101111000\\ 010011010111100\\ 011010111100010\\ 000000000000000\\ 010011010111100\\ 101111000100110\\ 111100010011010\\ 110101111000100\\ 011010111100010\\ 110101111000100\\ 101111000100110\\ 001001101011110\\ 001001101011110\\ 111100010011010\\ 000000000000000\\ 100110101111000\\ \end{array}\right].$
Figure 4: The autocorrelation magnitude distribution of $\mathbf{C}^{1,\alpha}$ in Example 2
Figure 5: The autocorrelation magnitude distribution of $\mathbf{C}^{1,\alpha^{2}}$ in Example 2
Figure 6: The correlation magnitude distribution of $\mathbf{C}^{1,\alpha}$ and $\mathbf{C}^{1,\alpha^{2}}$ in Example 2
III-B The second construction of periodic QCSSs from a special polynomial
In this subsection, we will use a special polynomial to construct periodic QCSSs with large set size.
Let $q=p^{n}$ with a prime $p$ and a positive integer $n$. Let $\alpha$ be a primitive element of ${\mathbb{F}}_{q}$. Assume that $d_{0},d_{1},\cdots,d_{q-1}$ are all the elements of ${\mathbb{F}}_{q}$, i.e. ${\mathbb{F}}_{q}=\{d_{0},d_{1},\cdots,d_{q-1}\}$. Let $f(x)=x^{p^{k}+1}+ax^{p^{k}}+bx+c\in{\mathbb{F}}_{q}[x]$, where $a,b,c\in{\mathbb{F}}_{q}$ and $\gcd(k,n)=1$. Let $\chi_{1}$ be the canonical additive character of ${\mathbb{F}}_{q}$. We define a constituent sequence $\mathbf{f}_{l}^{a,b,c}$ of period $q-1$ by
$\mathbf{f}_{l}^{a,b,c}=\left(\mathbf{f}_{l}^{a,b,c}(t)\right)_{t=0}^{q-2},\leavevmode\nobreak\ \leavevmode\nobreak\ \mbox{where }\mathbf{f}_{l}^{a,b,c}(t)=\chi_{1}(\alpha^{t}f(d_{l}))\mbox{ for }0\leq l\leq q-1.$ (14)
Then we obtain a two-dimensional $q\times(q-1)$ matrix
$\mathbf{F}^{a,b,c}=\left[\begin{array}[]{cccc}\mathbf{f}_{0}^{a,b,c}\\ \mathbf{f}_{1}^{a,b,c}\\ \vdots,\\ \mathbf{f}_{q-1}^{a,b,c}\end{array}\right].$
Such matrices form a complementary sequence set
=$\mathcal{F}=\left\{\mathbf{F}^{a,b,c}:a\in{\mathbb{F}}_{q},b\in{\mathbb{F}}_{q},c\in{\mathbb{F}}_{q}\right\}.$ (16)
Theorem III.3.
Let $p=2$ and $q=2^{n}$ for a positive integer $n\geq 3$. Let $f(x)=x^{p^{k}+1}+ax^{p^{k}}+bx+c\in{\mathbb{F}}_{q}[x]$, where $a,b,c\in{\mathbb{F}}_{q}$ and $\gcd(k,n)=1$. Let $\mathcal{F}$ be the complementary sequence set defined in (16). Then $\mathcal{F}$ is a periodic $(q^{3},q,q-1,2q)$-QCSS with alphabet size $2$ which is asymptotically near-optimal with respect to the correlation lower bound in (1).
Proof.
Let $p=2$, $q=2^{n}$ and $f(x)=x^{p^{k}+1}+ax^{p^{k}}+bx+c\in{\mathbb{F}}_{q}[x]$, where $a,b,c\in{\mathbb{F}}_{q}$ and $\gcd(k,n)=1$. For any two complementary sequences $\mathbf{F}^{a_{1},b_{1},c_{1}}$, $\mathbf{F}^{a_{2},b_{2},c_{2}}$ in $\mathcal{F}$ and $\tau\in[0,q-2]$, we have
$R_{\mathbf{F}^{a_{1},b_{1},c_{1}},\mathbf{F}^{a_{2},b_{2},c_{2}}}(\tau)$
$=$ $\sum_{l=0}^{q-1}R_{\mathbf{f}_{l}^{a_{1},b_{1},c_{1}},\mathbf{f}_{l}^{a_{2},b_{2},c_{2}}}(\tau)$
$=$ $\sum_{l=0}^{q-1}\sum_{t=0}^{q-2}\chi_{1}(\alpha^{t}(d_{l}^{p^{k}+1}+a_{1}d_{l}^{p^{k}}+b_{1}d_{l}+c_{1}))\overline{\chi_{1}}(\alpha^{t+\tau}(d_{l}^{p^{k}+1}+a_{2}d_{l}^{p^{k}}+b_{2}d_{l}+c_{2}))$
$=$ $\sum_{l=0}^{q-1}\sum_{t=0}^{q-2}\chi_{1}\left((1-\alpha^{\tau})\alpha^{t}d_{l}^{p^{k}+1}+(a_{1}-a_{2}\alpha^{\tau})\alpha^{t}d_{l}^{p^{k}}+(b_{1}-b_{2}\alpha^{\tau})\alpha^{t}d_{l}+(c_{1}-c_{2}\alpha^{\tau})\alpha^{t}\right),$
where $a_{1},b_{1},c_{1},a_{2},b_{2},c_{2}\in{\mathbb{F}}_{q}$.
We divide into the following cases to determine the value distribution of $R_{\mathbf{F}^{a_{1},b_{1},c_{1}},\mathbf{F}^{a_{2},b_{2},c_{2}}}(\tau)$.
Case 1: If $\tau=0$, then
$R_{\mathbf{F}^{a_{1},b_{1},c_{1}},\mathbf{F}^{a_{2},b_{2},c_{2}}}(\tau)$ (17)
$=$ $\sum_{l=0}^{q-1}\sum_{t=0}^{q-2}\chi_{1}\left((a_{1}-a_{2})\alpha^{t}d_{l}^{p^{k}}+(b_{1}-b_{2})\alpha^{t}d_{l}+(c_{1}-c_{2})\alpha^{t}\right)$
$=$ $\sum_{x\in{\mathbb{F}}_{q}}\sum_{y\in{\mathbb{F}}_{q}^{*}}\chi_{1}\left((a_{1}-a_{2})yx^{p^{k}}+(b_{1}-b_{2})yx+(c_{1}-c_{2})y\right).$
Subcase 1.1: If $a_{1}-a_{2}=0$ in Equation (17), by the orthogonal relation of the additive characters, we have
$R_{\mathbf{F}^{a_{1},b_{1},c_{1}},\mathbf{F}^{a_{2},b_{2},c_{2}}}(\tau)$
$=$ $\sum_{x\in{\mathbb{F}}_{q}}\sum_{y\in{\mathbb{F}}_{q}^{*}}\chi_{1}\left((b_{1}-b_{2})yx+(c_{1}-c_{2})y\right)$
$=$ $\left\{\begin{array}[]{ll}-q&\text{if}\leavevmode\nobreak\ b_{1}=b_{2}\leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ c_{1}\neq c_{2},\\ 0&\text{if}\leavevmode\nobreak\ b_{1}\neq b_{2}\leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ c_{1}=c_{2},\\ 0&\text{if}\leavevmode\nobreak\ b_{1}\neq b_{2}\leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ c_{1}\neq c_{2}.\end{array}\right.$
Subcase 1.2: If $a_{1}-a_{2}\neq 0$ in Equation (17), then
$R_{\mathbf{F}^{a_{1},b_{1},c_{1}},\mathbf{F}^{a_{2},b_{2},c_{2}}}(\tau)$
$=$ $\sum_{x\in{\mathbb{F}}_{q}}\sum_{y\in{\mathbb{F}}_{q}^{*}}\chi_{1}\left((a_{1}-a_{2})yx^{p^{k}}+(b_{1}-b_{2})yx+(c_{1}-c_{2})y\right)$
$=$ $\sum_{x\in{\mathbb{F}}_{q}}\sum_{y\in{\mathbb{F}}_{q}^{*}}\chi_{1}\left(\left((a_{1}-a_{2})x^{p^{k}}+(b_{1}-b_{2})x+(c_{1}-c_{2})\right)y\right).$
Now we consider the zeros of polynomial $g(x)=(a_{1}-a_{2})x^{p^{k}}+(b_{1}-b_{2})x+c_{1}-c_{2}$ in ${\mathbb{F}}_{2^{n}}$ with $a_{1}-a_{2}\neq 0$, where $p=2$ and $\gcd(n,k)=1$. Let $N_{g}$ denote the number of roots of $g(x)=0$ in ${\mathbb{F}}_{q}$. According to Lemma II.7, we know that $N_{g}\in\{0,1,2\}$ if $p=2$ and $\gcd(n,k)=1$. By the orthogonal relation of the additive characters, we have
$R_{\mathbf{F}^{a_{1},b_{1},c_{1}},\mathbf{F}^{a_{2},b_{2},c_{2}}}(\tau)$ (22)
$=$ $\left\{\begin{array}[]{ll}-q&\text{if}\leavevmode\nobreak\ N_{g}=0,\\ 0&\text{if}\leavevmode\nobreak\ N_{g}=1,\\ q&\text{if}\leavevmode\nobreak\ N_{g}=2.\end{array}\right.$
Case 2: If $\tau\neq 0$, then $1-\alpha^{\tau}\neq 0$. Thus we have
$R_{\mathbf{F}^{a_{1},b_{1},c_{1}},\mathbf{F}^{a_{2},b_{2},c_{2}}}(\tau)$
$=$ $\sum_{x\in{\mathbb{F}}_{q}}\sum_{y\in{\mathbb{F}}_{q}^{*}}\chi_{1}\left((1-\alpha^{\tau})yx^{p^{k}+1}+(a_{1}-a_{2}\alpha^{\tau})yx^{p^{k}}+(b_{1}-b_{2}\alpha^{\tau})yx+(c_{1}-c_{2}\alpha^{\tau})y\right)$
$=$ $\sum_{x\in{\mathbb{F}}_{q}}\sum_{y\in{\mathbb{F}}_{q}^{*}}\chi_{1}\left(\left((1-\alpha^{\tau})x^{p^{k}+1}+(a_{1}-a_{2}\alpha^{\tau})x^{p^{k}}+(b_{1}-b_{2}\alpha^{\tau})x+(c_{1}-c_{2}\alpha^{\tau})\right)y\right).$
Now we consider the polynomial $f(x)=(1-\alpha^{\tau})x^{p^{k}+1}+(a_{1}-a_{2}\alpha^{\tau})x^{p^{k}}+(b_{1}-b_{2}\alpha^{\tau})x+c_{1}-c_{2}\alpha^{\tau}$ in ${\mathbb{F}}_{2^{n}}$ with $1-\alpha^{\tau}\neq 0$, where $p=2$ and $\gcd(n,k)=1$. Let $N_{f}$ denote the number of zeros of $f$ in ${\mathbb{F}}_{q}$. According to Corollary II.6, we know that $N_{f}\in\{0,1,2,3\}$. By the orthogonal relation of additive characters, we then have
$R_{\mathbf{F}^{a_{1},b_{1},c_{1}},\mathbf{F}^{a_{2},b_{2},c_{2}}}(\tau)$ (27)
$=$ $\left\{\begin{array}[]{ll}-q&\text{if}\leavevmode\nobreak\ N_{f}=0,\\ 0&\text{if}\leavevmode\nobreak\ N_{f}=1,\\ q&\text{if}\leavevmode\nobreak\ N_{f}=2,\\ 2q&\text{if}\leavevmode\nobreak\ N_{f}=3.\end{array}\right.$
When $(a_{1},b_{1},c_{1},a_{2},b_{2},c_{2})$ runs through ${\mathbb{F}}_{q}^{6}$, each of the coefficients $a_{1}-a_{2}\alpha^{\tau}$, $b_{1}-b_{2}\alpha^{\tau}$, $a_{1}-a_{2}\alpha^{\tau}$ of $f(x)$ also runs through ${\mathbb{F}}_{q}$. By Corollary II.6, there exists $(a_{1},b_{1},c_{1},a_{2},b_{2},c_{2})\in{\mathbb{F}}_{q}^{6}$ such that $N_{f}=3$.
Based on all these cases discussed above, the maximum periodic correlation magnitude of $\mathcal{F}$ is $2q$. Thus $\mathcal{F}$ is a periodic $(q^{3},q,q-1,2q)$-QCSS. According to the correlation lower bound in (1), we have
$\vartheta_{{\text{opt}}}=q(q-1)\sqrt{\frac{q^{2}-1}{q^{3}(q-1)-1}}=q\sqrt{\frac{(q-1)^{3}(q+1)}{q^{3}(q-1)-1}}.$
It is easy to see that
$\lim_{q\rightarrow+\infty}\frac{\vartheta_{\max}}{\vartheta_{{\text{opt}}}}=\lim_{q\rightarrow+\infty}\frac{2q}{q\sqrt{\frac{(q-1)^{3}(q+1)}{q^{3}(q-1)-1)}}}=2.$
Then $\mathcal{F}$ is asymptotically near-optimal with respect to the correlation lower bound in (1). This completes the proof of this theorem.
Example 3.
Let $p=2$ and $n=3$. Then the parameters of the QCSS $\mathcal{F}$ constructed in Theorem III.3 are $(512,8,7,16)$ and its alphabet is given by $\{(-1)^{i}:i\in[0,1]\}$. By Magma program, we verify that the periodic correlation function $R_{\mathbf{F}^{a_{1},b_{1},c_{1}},\mathbf{F}^{a_{2},b_{2},c_{1}}}(\tau)$ is equal to 00, $8$ or $16$ for any $a,b,c\in{\mathbb{F}}_{q}$ and $0\leq\tau\leq 6$ except the trivial case that $a_{1}=a_{2},b_{1}=b_{2},c_{1}=c_{2}$ and $\tau=0$. For instance, the matrices $\mathbf{F}^{1,\alpha,\alpha}$ and $\mathbf{F}^{1,\alpha,\alpha^{2}}$ in $\mathcal{F}$ are presented as follows, where each entry stands for a power of $\zeta_{2}=-1$. By Python program, we show the autocorrelation magnitude distribution of $\mathbf{F}^{1,\alpha,\alpha}$ in Fig. 7, the autocorrelation magnitude distribution of $\mathbf{F}^{1,\alpha,\alpha^{2}}$ in Fig. 8, and the correlation magnitude distribution of $\mathbf{F}^{1,\alpha,\alpha}$ and $\mathbf{F}^{1,\alpha,\alpha^{2}}$ in Fig. 9, respectively.
$\mathbf{F}^{1,\alpha,\alpha}=\left[\begin{array}[]{cccc}0000000\\ 0010111\\ 0101110\\ 1001011\\ 0000000\\ 1110010\\ 0101110\\ 0101110\\ \end{array}\right],\ \mathbf{F}^{1,\alpha,\alpha^{2}}=\left[\begin{array}[]{cccc}1110010\\ 1100101\\ 1011100\\ 0111001\\ 1110010\\ 0000000\\ 1011100\\ 1011100\\ \end{array}\right].$
Figure 7: The autocorrelation magnitude distribution of $\mathbf{F}^{1,\alpha,\alpha}$ in Example 3
Figure 8: The autocorrelation magnitude distribution of $\mathbf{F}^{1,\alpha,\alpha^{2}}$ in Example 3
Figure 9: The correlation magnitude distribution of $\mathbf{F}^{1,\alpha,\alpha}$ and $\mathbf{F}^{1,\alpha,\alpha^{2}}$ in Example 3
III-C The third construction of periodic QCSSs from a cubic polynomial
In this subsection, we use a cubic polynomial to construct periodic QCSSs with large set size. Let $q=p^{n}$ with an odd prime $p$ and a positive integer $n$. Let $\alpha$ be a primitive element of ${\mathbb{F}}_{q}$. Assume that $d_{0},d_{1},\ldots,d_{q-1}$ are all the elements of ${\mathbb{F}}_{q}$, i.e. ${\mathbb{F}}_{q}=\{d_{0},d_{1},\ldots,d_{q-1}\}$. Let $f(x)=x^{3}+ax^{2}+bx+c\in{\mathbb{F}}_{q}[x]$, where $a,b,c\in{\mathbb{F}}_{q}$. Let $\chi_{1}$ be the canonical additive character of ${\mathbb{F}}_{q}$. We define a constituent sequence $\mathbf{h}_{l}^{a,b,c}$ of period $q-1$ by
$\mathbf{h}_{l}^{a,b,c}=\left(\mathbf{h}_{l}^{a,b,c}(t)\right)_{t=0}^{q-2},\leavevmode\nobreak\ \leavevmode\nobreak\ \text{where}\ \mathbf{h}_{l}^{a,b,c}(t)=\chi_{1}\left(\alpha^{t}f(d_{l})\right)\mbox{ for }0\leq l\leq q-1.$
Then we obtain a two-dimensional $q\times(q-1)$ matrix
$\mathbf{H}^{a,b,c}=\left[\begin{array}[]{cccc}\mathbf{h}_{0}^{a,b,c}\\ \mathbf{h}_{l}^{a,b,c}\\ \vdots\\ \mathbf{h}_{q-1}^{a,b,c}\\ \end{array}\right].$
Such matrices constitute a complementary sequence set
=$\mathcal{H}=\left\{\mathbf{H}^{a,b,c}:a\in{\mathbb{F}}_{q},b\in{\mathbb{F}}_{q},c\in{\mathbb{F}}_{q}\right\}.$ (30)
Theorem III.4.
Let $q=p^{n}$ with an odd prime $p$ and a positive integer $n$. Let $f(x)=x^{3}+ax^{2}+bx+c\in{\mathbb{F}}_{q}[x]$, where $a,b,c\in{\mathbb{F}}_{q}$. Let $\mathcal{H}$ be the complementary sequence set defined in (30). Then $\mathcal{H}$ is a periodic $(q^{3},q,q-1,2q)$-QCSS with alphabet size $p$ which is asymptotically near-optimal with respect to the correlation lower bound in(1).
Proof.
For any two complementary sequences $\mathbf{H}^{a_{1},b_{1},c_{1}}$, $\mathbf{H}^{a_{2},b_{2},c_{2}}$ of $\mathcal{H}$ and $\tau\in[0,q-2]$, we have
$R_{\mathbf{H}^{a_{1},b_{1},c_{1}},\mathbf{H}^{a_{2},b_{2},c_{2}}}(\tau)$
$=$ $\sum_{l=0}^{q-1}R_{\mathbf{h}_{l}^{a_{1},b_{1},c_{1}},\mathbf{h}_{l}^{a_{2},b_{2},c_{2}}}(\tau)$
$=$ $\sum_{l=0}^{q-1}\sum_{t=0}^{q-2}\chi_{1}\left(\alpha^{t}(d_{l}^{3}+a_{1}d_{l}^{2}+b_{1}d_{l}+c_{1})\right)\overline{\chi_{1}}\left(\alpha^{t+\tau}(d_{l}^{3}+a_{2}d_{l}^{2}+b_{2}d_{l}+c_{2})\right)$
$=$ $\sum_{l=0}^{q-1}\sum_{t=0}^{q-2}\chi_{1}\left((1-\alpha^{\tau})\alpha^{t}d_{l}^{3}+(a_{1}-a_{2}\alpha^{\tau})\alpha^{t}d_{l}^{2}+(b_{1}-b_{2}\alpha^{\tau})\alpha^{t}d_{l}+(c_{1}-c_{2}\alpha^{\tau})\alpha^{t}\right),$
where $a_{1},a_{2},b_{1},b_{2},c_{1},c_{2}\in{\mathbb{F}}_{q}$. We consider the following cases to determine its value distribution.
Case 1: If $\tau=0$, then
$R_{\mathbf{H}^{a_{1},b_{1},c_{1}},\mathbf{H}^{a_{2},b_{2},c_{2}}}(\tau)$ (31)
$=$ $\sum_{l=0}^{q-1}\sum_{t=0}^{q-2}\chi_{1}\left((a_{1}-a_{2})\alpha^{t}d_{l}^{2}+(b_{1}-b_{2})\alpha^{t}d_{l}+(c_{1}-c_{2})\alpha^{t}\right)$
$=$ $\sum_{x\in{\mathbb{F}}_{q}}\sum_{y\in{\mathbb{F}}_{q}^{*}}\chi_{1}\left((a_{1}-a_{2})yx^{2}+(b_{1}-b_{2})yx+(c_{1}-c_{2})y\right).$
Subcase 1.1: If $a_{1}-a_{2}=0$ in Equation (31), by the orthogonal relation of the additive characters, we have
$R_{\mathbf{H}^{a_{1},b_{1},c_{1}},\mathbf{H}^{a_{2},b_{2},c_{2}}}(\tau)$
$=$ $\sum_{x\in{\mathbb{F}}_{q}}\sum_{y\in{\mathbb{F}}_{q}^{*}}\chi_{1}((b_{1}-b_{2})yx+(c_{1}-c_{2})y)$
$=$ $\left\{\begin{array}[]{ll}-q&\mbox{if $b_{1}=b_{2}$ and $c_{1}\neq c_{2}$},\\ 0&\mbox{if $b_{1}\neq b_{2}$ and $c_{1}=c_{2}$},\\ 0&\mbox{if $b_{1}\neq b_{2}$ and $c_{1}\neq c_{2}$}.\end{array}\right.$
Subcase 1.2: If $a_{1}-a_{2}\neq 0$, then
$R_{\mathbf{H}^{a_{1},b_{1},c_{1}},\mathbf{H}^{a_{2},b_{2},c_{2}}}(\tau)$
$=$ $\sum_{x\in{\mathbb{F}}_{q}}\sum_{y\in{\mathbb{F}}_{q}^{*}}\chi_{1}((a_{1}-a_{2})yx^{2}+(b_{1}-b_{2})yx+(c_{1}-c_{2})y)$
$=$ $\sum_{x\in{\mathbb{F}}_{q}}\sum_{y\in{\mathbb{F}}_{q}^{*}}\chi_{1}\left(\left((a_{1}-a_{2})x^{2}+(b_{1}-b_{2})x+c_{1}-c_{2}\right)y\right).$
Now we consider the zeros of the polynomial $g(x)=(a_{1}-a_{2})x^{2}+(b_{1}-b_{2})x+c_{1}-c_{2}$ in ${\mathbb{F}}_{q}$ with $a_{1}-a_{2}\neq 0$, where $p$ is an odd prime. Let $N_{g}$ denote the number of roots of $g(x)=0$ in ${\mathbb{F}}_{q}$. It is easy to know that $N_{g}\in\{0,1,2\}$. By the orthogonal relation of the additive characters, we have
$R_{\mathbf{H}^{a_{1},b_{1},c_{1}},\mathbf{H}^{a_{2},b_{2},c_{2}}}(\tau)$ (36)
$=$ $\left\{\begin{array}[]{ll}-q&\mbox{if $N_{f}=0$},\\ 0&\mbox{if $N_{f}=1$},\\ q&\mbox{if $N_{f}=2$}.\end{array}\right.$
Case 2: If $\tau\neq 0$, then $1-\alpha^{\tau}\neq 0$. Thus we have
$R_{\mathbf{H}^{a_{1},b_{1},c_{1}},\mathbf{H}^{a_{2},b_{2},c_{2}}}(\tau)$
$=$ $\sum_{x\in{\mathbb{F}}_{q}}\sum_{y\in{\mathbb{F}}_{q}^{*}}\chi_{1}\left((1-\alpha^{\tau})yx^{3}+(a_{1}-a_{2}\alpha^{\tau})yx^{2}+(b_{1}-b_{2}\alpha^{\tau})yx+(c_{1}-c_{2}\alpha^{\tau})y\right)$
$=$ $\sum_{x\in{\mathbb{F}}_{q}}\sum_{y\in{\mathbb{F}}_{q}^{*}}\chi_{1}\left(\left((1-\alpha^{\tau})x^{3}+(a_{1}-a_{2}\alpha^{\tau})x^{2}+(b_{1}-b_{2}\alpha^{\tau})x+c_{1}-c_{2}\alpha^{\tau}\right)y\right).$
Now we consider the zeros of the polynomial $f(x)=(1-\alpha^{\tau})x^{3}+(a_{1}-a_{2}\alpha^{\tau})x^{2}+(b_{1}-b_{2}\alpha^{\tau})x+c_{1}-c_{2}\alpha^{\tau}$ in ${\mathbb{F}}_{q}$ with $1-\alpha^{\tau}\neq 0$, where $p$ is an odd prime. Let $N_{f}$ denote the number of zeros of $f$ in ${\mathbb{F}}_{q}$. It is known that $N_{f}\in\{0,1,2,3\}$. Note that the each of the coefficients $a_{1}-a_{2}\alpha^{\tau}$, $b_{1}-b_{2}\alpha^{\tau}$ and $c_{1}-c_{2}\alpha^{\tau}$ runs over ${\mathbb{F}}_{q}$ when all of $a_{1},a_{1},b_{1},b_{2},c_{1},c_{2}$ run over ${\mathbb{F}}_{q}$. Note that there exists a polynomial $f(x)=(1-\alpha^{\tau})(x-x_{1})(x-x_{2})(x-x_{3})$ for pairwise distinct elements $x_{1},x_{2},x_{3}\in{\mathbb{F}}_{q}$ which has three zeros in ${\mathbb{F}}_{q}$. By the orthogonal relation of additive characters, we then have
$R_{\mathbf{H}^{a_{1},b_{1},c_{1}},\mathbf{H}^{a_{2},b_{2},c_{2}}}(\tau)$ (41)
$=$ $\left\{\begin{array}[]{ll}-q&\mbox{if $N_{f}=0$},\\ 0&\mbox{if $N_{f}=1$},\\ q&\mbox{if $N_{f}=2$},\\ 2q&\mbox{if $N_{f}=3$}.\end{array}\right.$
Based on all the cases discussed above, the maximum periodic correlation magnitude of $\mathcal{H}$ is $2q$. Thus $\mathcal{H}$ is a periodic $(q^{3},q,q-1,2q)$-QCSS. According to the correlation lower bound in (1), we have
$\vartheta_{{\text{opt}}}=q(q-1)\sqrt{\frac{q^{2}-1}{q^{3}(q-1)-1}}=q\sqrt{\frac{(q-1)^{3}(q+1)}{q^{3}(q-1)-1}}.$
It is easy to see that
$\lim_{q\rightarrow+\infty}\frac{\vartheta_{\max}}{\vartheta_{{\text{opt}}}}=\lim_{q\rightarrow+\infty}\frac{2q}{q\sqrt{\frac{(q-1)^{3}(q+1)}{q^{3}(q-1)-1)}}}=2.$
Then $\mathcal{H}$ is asymptotically near-optimal with respect to the correlation lower bound in (1). This completes the proof of this theorem.
Example 4.
Let $p=3$ and $n=2$. Then the parameters of the QCSS $\mathcal{H}$ constructed in Theorem III.4 are $(729,9,8,18)$ and its alphabet is given by $\{e^{2\pi\sqrt{-1}i/3}:i\in[0,2]\}$. By Magma program, we verify that the periodic correlation function $R_{\mathbf{H}^{a_{1},b_{1},c_{1}},\mathbf{H}^{a_{2},b_{2},c_{1}}}(\tau)$ is equal to 00, $9$ or $18$ for any $a,b,c\in{\mathbb{F}}_{q}$ and $0\leq\tau\leq 7$ except the trivial case that $a_{1}=a_{2},b_{1}=b_{2},c_{1}=c_{2}$ and $\tau=0$. For instance, the matrices $\mathbf{H}^{1,\alpha,\alpha}$ and $\mathbf{H}^{1,\alpha,\alpha^{2}}$ in $\mathcal{H}$ are presented as follows, where each entry stands for a power of $\zeta_{3}=e^{2\pi\sqrt{-1}/3}$. By Python program, we show the autocorrelation magnitude distribution of $\mathbf{H}^{1,\alpha,\alpha}$ in Fig. $10$, the autocorrelation magnitude distribution of $\mathbf{H}^{1,\alpha,\alpha^{2}}$ in Fig. $11$, and the correlation magnitude distribution of $\mathbf{H}^{1,\alpha,\alpha}$ and $\mathbf{H}^{1,\alpha,\alpha^{2}}$ in Fig. $12$. We remark that the upper bound of $\vartheta_{\max}$ in Theorem III.4 is tight in this case.
$\mathbf{h}^{1,\alpha,\alpha}=\left[\begin{array}[]{cccc}22101120\\ 02210112\\ 22101120\\ 11202210\\ 00000000\\ 22101120\\ 20221011\\ 02210112\\ 01120221\\ \end{array}\right],\ \mathbf{h}^{1,\alpha,\alpha^{2}}=\left[\begin{array}[]{cccc}02210112\\ 12022101\\ 02210112\\ 21011202\\ 10112022\\ 02210112\\ 00000000\\ 12022101\\ 11202210\\ \end{array}\right].$
Figure 10: The autocorrelation magnitude distribution of $\mathbf{H}^{1,\alpha,\alpha}$ in Example 4
Figure 11: The autocorrelation magnitude distribution of $\mathbf{H}^{1,\alpha,\alpha^{2}}$ in Example 4
Figure 12: The correlation magnitude distribution of $\mathbf{H}^{1,\alpha,\alpha}$ and $\mathbf{H}^{1,\alpha,\alpha^{2}}$ in Example 4
III-D The fourth construction of periodic QCSSs from Gaussian sums
Let $q=p^{n}$ with a prime $p$ and a positive integer $n$. Let $\alpha$ be a primitive element of ${\mathbb{F}}_{q}$. Assume that $d_{0},d_{1},\cdots,d_{q-2}$ are all the elements of ${\mathbb{F}}_{q}^{*}$, i.e. ${\mathbb{F}}_{q}^{*}=\{d_{0},d_{1},\cdots,d_{q-2}\}$. Let $\chi_{b}$, $b\in{\mathbb{F}}_{q}$, and $\varphi_{i}$, $20\leq i\leq q-$ 2, be additive character and multiplicative characters of ${\mathbb{F}}_{q}$ defined in Subsection II-B, respectively.
For each $b\in{\mathbb{F}}_{q}$ and $20\leq i\leq q-$ 2, we define a constituent sequence $\mathbf{s}_{l}^{i,b}$ of period $q-1$ as
$\mathbf{s}_{l}^{i,b}=\left(\mathbf{s}_{l}^{i,b}(t)\right)_{t=0}^{q-2},\leavevmode\nobreak\ \text{where}\ \mathbf{s}_{l}^{i,b}=\varphi_{i}(d_{l})\chi_{b}(\alpha^{t}d_{l})\chi_{1}({\alpha^{t}})\leavevmode\nobreak\ \text{for}\ 0\leq l\leq q-2.$
Then we obtain a two-dimensional $(q-1)\times(q-1)$ matrix
$\mathbf{S}^{i,b}=\left[\begin{array}[]{cccc}\mathbf{s}_{0}^{i,b}\\ \mathbf{s}_{1}^{i,b}\\ \vdots,\\ \mathbf{s}_{q-2}^{i,b}\end{array}\right].$
The set
$\mathcal{S}=\left\{\mathbf{S}^{i,b}:0\leq i\leq q-2,b\in{\mathbb{F}}_{q}\right\}$ (44)
is a complementary sequence set.
Theorem III.5.
Let $q=p^{n}$ with a prime $p$ and a positive integer $n$. Let $\mathcal{S}$ be the complementary sequence set defined in (44). Then $\mathcal{S}$ is a periodic $(q^{2}-q,q-1,q-1,q)$-QCSS with alphabet size $p(p^{n}-1)$ which is asymptotically optimal with respect to the correlation lower bound in (1).
Proof.
For any two complementary sequences $\mathbf{S}^{i_{1},b_{1}}$, $\mathbf{S}^{i_{2},b_{2}}$ in $\mathcal{S}$ and $\tau\in[0,q-2]$, we have
$R_{\mathbf{S}^{i_{1},b_{1}},\mathbf{S}^{i_{2},b_{2}}}(\tau)$
$=$ $\sum_{l=0}^{q-2}R_{\mathbf{s}_{l}^{i_{1},b_{1}},\mathbf{s}_{l}^{i_{2},b_{2}}}(\tau)$
$=$ $\sum_{l=0}^{q-2}\sum_{t=0}^{q-2}\varphi_{i_{1}}(d_{l})\chi_{b_{1}}(d_{l}\alpha^{t})\chi_{1}(\alpha^{t})\overline{\varphi_{i_{2}}}(d_{l})\overline{\chi_{b_{2}}}(d_{l}\alpha^{t+\tau})\overline{\chi_{1}}(\alpha^{t+\tau})$
$=$ $\sum_{l=0}^{q-2}\sum_{t=0}^{q-2}\varphi_{i_{1}-i_{2}}(d_{l})\chi_{1}\left((b_{1}-b_{2}\alpha^{\tau})\alpha^{t}d_{l}\right)\chi_{1}\left((1-\alpha^{\tau})\alpha^{t}\right),$
where $0\leq i_{1}\leq q-2,0\leq i_{2}\leq q-2,b_{1},b_{2}\in{\mathbb{F}}_{q}$. Then we consider the following cases to determine the value distribution of $R_{\mathbf{S}^{i_{1},b_{1}},\mathbf{S}^{i_{2},b_{2}}}(\tau)$.
Case 1: If $\tau=0$, $i_{1}=i_{2}$ and $b_{1}\neq b_{2}$, then
$R_{\mathbf{S}^{i_{1},b_{1}},\mathbf{S}^{i_{2},b_{2}}}(\tau)$
$=$ $\sum_{l=0}^{q-2}\sum_{t=0}^{q-2}\chi_{1}\left((b_{1}-b_{2})\alpha^{t}d_{l}\right)$
$=$ $\sum_{l=0}^{q-2}\sum_{x\in{\mathbb{F}}_{q}^{*}}\chi_{1}\left((b_{1}-b_{2})d_{l}x\right).$
By the orthogonal relation of the additive characters, we have
$R_{\mathbf{S}^{i_{1},b_{1}},\mathbf{S}^{i_{2},b_{2}}}(\tau)=-(q-1).$
Case 2: If $\tau=0$, $i_{1}\neq i_{2}$ and $b_{1}=b_{2}$, then
$R_{\mathbf{S}^{i_{1},b_{1}},\mathbf{S}^{i_{2},b_{2}}}(\tau)$
$=$ $\sum_{l=0}^{q-2}\sum_{t=0}^{q-2}\varphi_{i_{1}-i_{2}}(d_{l})$
$=$ $(q-1)\sum_{y\in{\mathbb{F}}_{q}^{*}}\varphi_{i_{1}-i_{2}}(y)=0$
by the orthogonal relation of the multiplicative characters.
Case 3: If $\tau=0$, $i_{1}\neq i_{2}$ and $b_{1}\neq b_{2}$, then
$R_{\mathbf{S}^{i_{1},b_{1}},\mathbf{S}^{i_{2},b_{2}}}(\tau)$
$=$ $\sum_{l=0}^{q-2}\sum_{t=0}^{q-2}\varphi_{i_{1}-i_{2}}(d_{l})\chi_{1}((b_{1}-b_{2})\alpha^{t}d_{l})$
$=$ $\sum_{y\in{\mathbb{F}}_{q}^{*}}\varphi_{i_{1}-i_{2}}(y)\sum_{x\in{\mathbb{F}}_{q}^{*}}\chi_{1}((b_{1}-b_{2})yx)$
$=$ $-\sum_{y\in{\mathbb{F}}_{q}^{*}}\varphi_{i_{1}-i_{2}}(y)=0$
by the orthogonal relation of the additive and multiplicative characters.
Case 4: If $\tau\neq 0$ and $i_{1}=i_{2}$, then
$R_{\mathbf{S}^{i_{1},b_{1}},\mathbf{S}^{i_{2},b_{2}}}(\tau)$
$=$ $\sum_{l=0}^{q-2}\sum_{t=0}^{q-2}\chi_{1}((b_{1}-b_{2}\alpha^{\tau})\alpha^{t}d_{l})\chi_{1}((1-\alpha^{\tau})\alpha^{t})$
$=$ $\sum_{x\in{\mathbb{F}}_{q}^{*}}\chi_{1}((1-\alpha^{\tau})x)\sum_{y\in{\mathbb{F}}_{q}^{*}}\chi_{1}((b_{1}-b_{2}\alpha^{\tau})xy)$
$=$ $\left\{\begin{array}[]{ll}(q-1)\sum_{x\in{\mathbb{F}}_{q}^{*}}\chi_{1}((1-\alpha^{\tau})x)&\text{if}\leavevmode\nobreak\ b_{1}-b_{2}\alpha^{\tau}=0\\ -\sum_{x\in{\mathbb{F}}_{q}^{*}}\chi_{1}((1-\alpha^{\tau})x)&\text{if}\leavevmode\nobreak\ b_{1}-b_{2}\alpha^{\tau}\neq 0\end{array}\right.$
$=$ $\left\{\begin{array}[]{ll}-(q-1)&\text{if}\leavevmode\nobreak\ b_{1}-b_{2}\alpha^{\tau}=0,\\ 1&\text{if}\leavevmode\nobreak\ b_{1}-b_{2}\alpha^{\tau}\neq 0,\end{array}\right.$
according to the orthogonal relation of the additive characters.
Case 5: If $\tau\neq 0$ and $i_{1}\neq i_{2}$, then
$R_{\mathbf{S}^{i_{1},b_{1}},\mathbf{S}^{i_{2},b_{2}}}(\tau)$
$=$ $\sum_{l=0}^{q-2}\sum_{t=0}^{q-2}\varphi_{i_{1}-i_{2}}(d_{l})\chi_{1}((b_{1}-b_{2}\alpha^{\tau})\alpha^{t}d_{l})\chi_{1}((1-\alpha^{\tau})\alpha^{t})$
$=$ $\sum_{x\in{\mathbb{F}}_{q}^{*}}\sum_{y\in{\mathbb{F}}_{q}^{*}}\varphi_{i_{1}-i_{2}}(y)\chi_{1}((b_{1}-b_{2}\alpha^{\tau})xy)\chi_{1}((1-\alpha^{\tau})x)$
$=$ $\left\{\begin{array}[]{ll}\sum\limits_{y\in{\mathbb{F}}_{q}^{*}}\varphi_{i_{1}-i_{2}}(y)\sum\limits_{x\in{\mathbb{F}}_{q}^{*}}\chi_{1}((1-\alpha^{\tau})x)&\text{if}\leavevmode\nobreak\ b_{1}-b_{2}\alpha^{\tau}=0\\ \begin{subarray}{c}\mbox{$\sum\limits_{x\in{\mathbb{F}}_{q}^{*}}\chi_{1}((1-\alpha^{\tau})x)\overline{\varphi_{i_{1}-i_{2}}}((b_{1}-b_{2}\alpha^{\tau})x)\cdot$}\\ \mbox{$\sum\limits_{y\in{\mathbb{F}}_{q}^{*}}\varphi_{i_{1}-i_{2}}((b_{1}-b_{2}\alpha^{\tau})xy)\chi_{1}((b_{1}-b_{2}\alpha^{\tau})xy)$}\end{subarray}&\text{if}\leavevmode\nobreak\ b_{1}-b_{2}\alpha^{\tau}\neq 0\end{array}\right.$
$=$ $\left\{\begin{array}[]{ll}0&\text{if}\leavevmode\nobreak\ b_{1}-b_{2}\alpha^{\tau}=0,\\ \overline{\varphi_{i_{1}-i_{2}}}(b_{1}-b_{2}\alpha^{\tau})\varphi_{i_{1}-i_{2}}(1-\alpha^{\tau})G(\varphi_{i_{1}-i_{2}},\chi_{1})G(\overline{\varphi_{i_{1}-i_{2}}},\chi_{1})&\text{if}\leavevmode\nobreak\ b_{1}-b_{2}\alpha^{\tau}\neq 0,\end{array}\right.$
by the orthogonal relation of the additive and multiplicative characters and the definition of Gaussian sums. It is known that $G(\overline{\varphi_{i_{1}-i_{2}}},\chi_{1})=\varphi_{i_{1}-i_{2}}(-1)\overline{G(\varphi_{i_{1}-i_{2}},\chi_{1})}$. Then
$G(\varphi_{i_{1}-i_{2}},\chi_{1})G(\overline{\varphi_{i_{1}-i_{2}}},\chi_{1})=\varphi_{i_{1}-i_{2}}(-1)|G(\varphi,\chi_{1})|^{2}=q\varphi_{i_{1}-i_{2}}(-1)$
by Lemma II.2. Thus
$|R_{\mathbf{C}^{a_{1},b_{1}},\mathbf{C}^{a_{2},b_{2}}}(\tau)|\in\{0,\pm q\}.$
Based on the above discussions, we deduce that the maximum periodic correlation magnitude of $\mathcal{S}$ is $q$. Then $\mathcal{S}$ is a periodic $(q^{2}-q,q-1,q-1,q)$-QCSS. According to the correlation bound in (1), we have
$\vartheta_{{\text{opt}}}=\sqrt{\frac{(q-1)^{5}}{q(q-1)^{2}-1}}.$
It is obvious that
$\lim_{q\rightarrow+\infty}\frac{\vartheta_{\max}}{\vartheta_{{\text{opt}}}}=\lim_{q\rightarrow+\infty}\frac{q}{\sqrt{\frac{(q-1)^{5}}{q(q-1)^{2}-1}}}=1.$
Then $\mathcal{S}$ is asymptotically optimal with respect to the correlation lower bound in (1). This completes the proof of this theorem. ∎
Example 5.
Let $p=2$ and $n=3$. Then the parameters of the QCSS $\mathcal{S}$ constructed in Theorem III.5 are $(56,7,7,8)$. By Magma program, the periodic correlation function $R_{\mathbf{S}^{i_{1},b_{1}},\mathbf{S}^{i_{2},b_{2}}}(\tau)$ is equal to 00, $1$, $7$ or $8$ for any $b\in{\mathbb{F}}_{q}$, $0\leq i\leq 6$ and $0\leq\tau\leq 6$ except the trivial case that $i_{1}=i_{2},b_{1}=b_{2}$ and $\tau=0$. By Python program, we show the autocorrelation magnitude distribution of $\mathbf{S}^{1,\alpha}$ in Fig. 13, the autocorrelation magnitude distribution of $\mathbf{S}^{1,\alpha^{2}}$ in Fig. 14, and the correlation magnitude distribution of $\mathbf{S}^{1,\alpha}$ and $\mathbf{S}^{1,\alpha^{2}}$ in Fig. 15, respectively. These correlation magnitude distributions coincide with the results in Theorem III.5.
Figure 13: The autocorrelation magnitude distribution of $\mathbf{S}^{1,\alpha}$ in Example 5
Figure 14: The autocorrelation magnitude distribution of $\mathbf{S}^{1,\alpha^{2}}$ in Example 5
Figure 15: The correlation magnitude distribution of $\mathbf{S}^{1,\alpha}$ and $\mathbf{S}^{1,\alpha^{2}}$ in Example 5
IV Concluding remarks
In this paper, we presented five new families of asymptotically optimal or near-optimal periodic QCSSs with large set sizes from some special polynomials and Gaussian sums. Specially, we constructed the first two families of periodic QCSSs with set size $\Theta(K^{3})$ and flock size $K$. Most of the QCSSs constructed in this paper have very small alphabet size. The parameters of the QCSSs in this paper improve those of known ones in [10, 22]. From the figures in this paper, some complementary sequences have zero nontrivial autocorrelation magnitude. These advantages indicate that our QCSSs have potential applications in MC-CDMA communication systems.
Note that the correlation lower bound in (1) may not be tight if $M$ is too larger than $K$. It is interesting to improve this bound. The reader is invited to study this interesting topic.
Declarations
Conflict of interest The authors declare that they have no conflicts of interest relevant to the content of this article.